With j1=7.6%, what is equivalent j12?
1
Expert's answer
2021-10-25T16:08:31-0400

j1=7.6%

An Interest rate compounded annually is equivalent to an Effective Annual RATE(EAR).

Calculating Nominal Interest rate compounded monthly(j12) using EAR formula:-

EAR"=(1+r\/m)^m-1"

Where, r = Nominal Interest rate

m = no of times compounding in a year = 12

"0.076=(1+r\/12)^{12}-1"

"1.076=(1+r\/12)^{12}"

"1.076^{1\/12}=1+r\/12"

"1.0061-1=r\/12"

"0.0061=r\/12"

"r=0.0732=7.32\%"

"j_{12}=7.32\%"